OpenStudy (daenio):

Matt and his wife have 20 good friends, twelve of them are male and eight of them are female. They decide to have a diner party but can invite only eight guests. They decide to invite their guests by drawing their names from a hat. What is the probability that a) There will be an equal number of males and females at the dinner party? b) Mandie will be among those invited? c) There will be only one female guest? d) They will have only male guests? e) There will be at least one female guest?

OpenStudy (anonymous):

a) "equal number of males and females" which means 4 each. 12C4 * 8C4 / 20C8 = 495*70/125970 = 1155/4199 = 0.275 b) Mandie will be among those invited and Mandie is a female, so she must be chosen, 19C7 / 20C8 = 50388/125970 = 2/5 = 0.4 c) "There will be only one female guest" which means 7 males, 1 female, 12C7 * 8C1 / 20C8 = 792*8/125970 = 6336/125970 = 1056/20995 = 0.05 d) "They will have only male guests" which means 1 male, 7 females, 12C1 * 8C7 / 20C8 = 12*8/125970 = 96/125970 = 0.000762 e) "There will be at least one female guest"..., you have to consider the reversed way, i.e., 1 - (when there is no female guest at all) 1 - (12C8 / 20C8) = 1 - (495/125970) = 1 - 0.00393 = 0.996